# arXiv:2603.01316v1[eess.AS] 01 Mar 2026

### v1

@@ III Proposed Two-Stage Method
In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$
$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}(\text{mix}),$
@@
The audio encoder consists of a self-supervised speech model (wav2vec2-large-xlsr-53222https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
@@
We define the logit as the difference between these similarity scores as
@@
which represents how much more similar the first separated signal $\hat{s}_{1}$ is to the text prompt compared with the second signal $\hat{s}_{2}$. The logit is then mapped to a probability using the sigmoid function:
$\hat{y}=\sigma(\text{logit}),$
where $\hat{y}\in[0,1]$ indicates the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the text prompt. Finally, the predicted target speaker index is determined as
@@
During inference, the signal with the higher similarity is selected as the target speech $\hat{s}_{\text{tar}}$.
@@ IV Evaluation
The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems, and (ii) examine whether relative cues yield better performance than independent cues. The following sections provide an overview of the experimental dataset and training configurations, followed by detailed descriptions of each experiment, including the setup, results, and analysis.
@@ IV-A Dataset
To ensure broad linguistic and acoustic diversity, audio mixtures were constructed from multiple publicly available corpora spanning five major languages—Chinese, English, French, German, and Spanish—each providing annotated attributes such as language, gender, emotion, age, and transcription (see Table II). Each mixture contains speech from two speakers—one designated as the target and the other as the interference—together with a text prompt describing their relative characteristics.
@@
(1) Extraction of Speech Attributes: To provide ground-truth labels for the relative cue categories in each mixture, a set of speech attributes is extracted for every utterance. These attributes serve as the basis for constructing inter-speaker relative cues (e.g., higher pitch, faster speaking rate). All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or by mixture-level acoustic artifacts.
@@
Mean $F_{0}$ and $F_{0}$ span are extracted using the Librosa pYIN library [^4] https://librosa.org/doc/main/generated/librosa.pyin.html. When computing speaking duration, word- and pause-level durations are first obtained. If transcription annotations are available, these durations are extracted from time boundaries provided by the Montreal Forced Aligner [^5] https://montreal-forced-aligner.readthedocs.io/en/latest/; otherwise, they are estimated using the Silero voice activity detection tool [^6] https://github.com/snakers4/silero-vad. Speaking duration is then defined as the sum of word durations and short pauses, excluding pauses longer than 0.6 s, which are treated as silence. Speaking rate is computed as the total number of syllables divided by the speaking duration. Syllable counts are estimated based on vowel sequences: in Chinese, each character is treated as one syllable, whereas in English, French, German, and Spanish, syllables are approximated by counting contiguous vowel groups within words. The RMS energy is computed from the reverberant speech while accounting for the speaking duration. The speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone. Language, gender, age, emotional state, and transcription content are obtained directly from the original dataset annotations.
@@
(2) Audio Mixtures: Each corpus is split into non-overlapping training, validation, and test subsets, with no overlap in speakers or spoken words. We generated 100,000 mixtures for training and 10,000 mixtures each for validation and testing. To create realistic speech mixtures, each mixture is composed of two speech signals, $S_{1}$ and $S_{2}$, from different speakers. Both signals are first processed to remove silence—using forced alignment for transcribed speech or voice activity detection for untranscribed segments—and then temporally arranged with partial overlaps to mimic conversational dynamics. If either $S_{1}$ or $S_{2}$ is shorter than 3 s, the overlap duration is set to the length of the shorter signal, and zero-padding is applied at the beginning of the shorter signal using a randomly selected start offset uniformly drawn from $0,\text{duration of longer sample}-\text{duration of shorter sample}]$. For longer signals, the overlap duration is set to $(\text{length}(S_{1})+\text{length}(S_{2})-6)$ s, with $S_{2}$ padded using a fixed start offset of $6-\text{length}(S_{2})$ s.
@@
To incorporate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library [9], simulating different speaker positions within the room. The simulated rooms vary in size, reverberation time, and speaker–microphone distance. Room dimensions are randomly sampled within 9–11 m in length and width, and 2.6–3.5 m in height, while reverberation times (RT60) are sampled from 0.3–0.6 s. The microphone is positioned at the center of the room, and speaker locations are randomly assigned such that horizontal speaker–microphone distances range from 0.3 to 1.5 m, with speaker heights between 1.6 and 1.9 m. We keep the reverberant $S_{2}$ fixed and adjust the reverberant $S_{1}$ by applying a signal-to-interference ratio (SIR), uniformly sampled from –6 to 6 dB. For each mixture, one of the reverberant signals, $S_{1}$ or $S_{2}$, is randomly selected to serve as the target speech.
TABLE III: Summary of used speech datasets across five languages, detailing the number of speakers in the training, validation, and test splits, along with their partial attributes.
Dataset Language Speakers Attributes
CASIA [60] Chinese 2/1/1 Language, Emotion, Gender, Transcription
ESD [63] Chinese 6/2/2 Language, Emotion, Gender, Transcription
Aishell3 [49] Chinese 120/30/30 Language, Gender, Transcription
MagicData-Read [36] Chinese 33/8/11 Language, Age, Gender
Librispeech [38] English 120/30/30 Language, Gender, Transcription
ESD [63] English 6/2/2 Language, Emotion, Gender, Transcription
OreauFR-02 [21] French 21/7/4 Language, Emotion, Age, Gender
MLS French [39] French 90/18/18 Language, Gender, Transcription
EmoDB [6] German 6/2/2 Language, Emotion, Age, Gender, Transcription
MLS German [39] German 120/30/30 Language, Gender, Transcription
MLS Spanish [39] Spanish 79/20/20 Language, Gender, Transcription
EmoMatchSpanish [14] Spanish 30/10/10 Language, Emotion, Gender, Transcription
@@
(3) Text Prompts: As detailed in Section II, Table IV summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power.
@@
TABLE IV: Sample counts of each relative cue in the training, validation and test sets
@@ IV-B Training Configurations
For the speech encoder used in the classification model of the proposed two-stage framework, all cue types employed during training span linguistic, paralinguistic, and acoustic information. Therefore, the speech encoder plays a central role in capturing these diverse representations. We adopt the encoder from the pretrained wav2vec2-large-xlsr-53 model, a multilingual self-supervised speech representation model based on the wav2vec 2.0 architecture [4]. Trained on speech data from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, wav2vec2-large-xlsr-53 learns universal, language-independent representations directly from raw audio. Compared to the state-of-the-art speaker encoder ECAPA-TDNN, which is primarily designed to extract speaker-specific acoustic embeddings, wav2vec2-large-xlsr-53 provides richer representations that jointly capture linguistic, paralinguistic, and acoustic information. This makes it particularly well suited for cross-lingual target speech extraction guided by diverse cue types. For wav2vec2-large-xlsr-53, only the first five encoder layers are used, with fine-tuning applied to the fifth layer.
@@
To encode text prompts into embeddings, we employ the LLaMA 3.2 1B Instruct model for all single-stage TSE systems as well as for the classification model in the two-stage framework. The LLaMA 3.2 1B instruction model was fine-tuned using LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
@@
To ensure high-quality signal separation, we adopt the SOTA speech separation model TF-Locoforme [43], which has demonstrated superior performance on public speech separation benchmarks and robustness under reverberant conditions. Both single-stage TSE models and the two-stage separation model are trained with the negative SI-SDR loss using float16 mixed precision and a batch size of 24, as is the two-stage classification model.
@@
All models were optimized using AdamW [29], with learning rates and gradient clipping configured individually for each model. The DPT-based single-stage TSE was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and linear warm-up over 1,000 updates. The TF-Locoformer-based single-stage TSE was trained with a learning rate of 1e-4 and a maximum norm of 5, whereas the two-stage separation model used a learning rate of 1e-3 with the same maximum norm. The classification models were trained with a learning rate of 1e-4 and a maximum norm of 30. Different learning rates and maximum norm values were adopted to account for the distinct architectures and convergence behaviors of the respective models. The learning rate was halved if the validation loss did not improve for three consecutive epochs, and training employed early stopping with a maximum of 100 epochs.
IV-C Two-Stage Method versus Single-Stage Baselines
Experimental Setup: We evaluate the proposed two-stage method against single-stage baselines. In the first stage, the separation model is trained from scratch to obtain separated speech signals. In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, where each prompt corresponds to one of the fourteen relative cue types listed in Table III. The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model. The latter is derived by adapting the TF-Locoformer speech separation model for target speech extraction. Specifically, rather than estimating two masks (one per speaker) to separate both sources, the model is reconfigured to predict a single time–frequency mask corresponding exclusively to the target speaker. Samples labeled as “Similar” and “Same” categories are excluded from the training, validation, and test sets due to their lack of discriminative power. Additionally, we train and evaluate a classification model that uses an enrollment utterance from the target speaker to enable comparison with textual relative cues.
Results and Analysis:
@@ IV-C Two-Stage Method versus Single-Stage Baselines
TABLE V: Comparison of single-stage and two-stage TSE systems across different cues
@@
Enrollment Audio 95.795.7% $15.2$ $3.31$
@@
We use scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [24] and perceptual evaluation of speech quality (PESQ) [42] to evaluate target speech extraction performance. For the TF-Locoformer-based two-stage TSE framework, we report three metrics: the separation performance of TF-Locoformer (Sep SI-SDRi/PESQ), the classification accuracy on the separated signals (Sep ACC), and the overall target speech extraction performance (SI-SDRi/PESQ). The overall performance is calculated as the mean score between the predicted and ground-truth target speech under various relative cue conditions. The results are summarized in Table V.
To evaluate performance for different cue types, we follow a prompt-based setup: Individual cues: Each template prompt contains only a single relative cue type. For each cue type, the model is evaluated on mixtures corresponding to that cue, producing cue-specific SI-SDRi, PESQ, and classification metrics. All cue: Each prompt includes all available single relative cues presented in the target speech of a mixture. The model is evaluated on mixtures with multiple cue types simultaneously, allowing us to assess performance under the combination of all available cues. Random cue: Each prompt contains a random subset of relative cues, and evaluation is performed on mixtures corresponding to these randomly selected cues. This measures the robustness of the two-stage method under partial or unpredictable cue guidance.
@@
The two-stage TSE also supports target speaker enrollment utterances as guidance, where the separated signal with the highest cosine similarity to the enrollment embedding is selected as the target. This enables a direct comparison with text-based relative cues. From the results, the TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based single-stage TSE, achieving significant 3+ dB improvements in SI-SDRi and 0.7–1.1 improvements in PESQ, except for the age category, which remains challenging for both models. This suggests that age prediction performance is bottlenecked by the cue discriminability rather than the model architecture. The classification performance of different cue types reflects their relative contribution to the single-stage TSE baselines. Compared with the single-stage TF-Locoformer-based TSE, the two-stage method demonstrates clear advantages. It achieves over 3 dB SI-SDRi improvement for the random cue, all cue, language, transcription, gender, and speaking rate cues, and gains more than 2 dB SI-SDRi for the remaining cue types (except for the age category cue). Additionally, it provides an average improvement of over 0.5 in PESQ across all cue types, highlighting its strong practical potential.
@@
Interestingly, some relative cue types—including all cue, random cue, language, gender, loudness, distance, temporal order and speaking duration—can even outperform the enrollment speech of the target speaker. This suggests that enrollment speech may contain potentially confounding information, as it could be recorded in a different environment and exhibit speech characteristics that differ from the target speech. These findings further highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging some distinguishable cue types present in the mixture speech samples.
@@
To establish an oracle benchmark, we additionally train and evaluate a separate classification model on the target–inference reverberant–clean speech pairs from each mixture described in Section IV (Audio Mixtures of Dataset). The evaluation results are shown in Table VI.
TABLE VI: Performance across different cues
Cue Type Clean ACC Sep ACC
Random Cue 99.3% 99.2%
All Cue 99.8% 99.8%
Language 99.5% 99.3%
Transcription 90.8% 90.4%
Gender 98.0% 98.1%
Emotion 85.8% 87.2%
Pitch Level 95.2% 94.6%
Pitch Range 88.4% 88.7%
Loudness 98.8% 97.2%
Distance 99.0% 98.4%
Age Category 64.0% 64.3%
Temporal Order 100% 99.8%
Speaking Rate 91.5% 91.3%
Speaking Duration 96.3% 95.9%
Enrollment Audio 96.1% 95.9%
In general, classification performance on clean speech (Clean ACC) generally exceeds that on separated speech (Sep ACC), as expected. Interestingly, the accuracy for each cue type closely matches that of the model trained on separated speech generated by the TF-Locoformer separation model, demonstrating the high quality of the separated signals. The varying classification accuracies across cue types also provide insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
Figure 4: Overall trend of classification accuracy with respect to the attribute differences between target and interfering speakers for each cue type.
We further analyze classification accuracy as a function of the continuous-valued attribute difference between paired target–interference clean speech samples for each cue type. The temporal order cue is excluded due to its perfect classification accuracy (100%). For the remaining cues, logistic regression is used to obtain a smooth estimate of overall classification accuracy over the continuous attribute space. As shown in Fig. 4, accuracy exhibits a clear increasing trend as attribute differences grow. Except for age category, all cues reach near-perfect classification when the differences become sufficiently large.

### v2

@@ III Proposed Two-Stage Method
In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$ as
$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}({mix}),$
@@
The audio encoder consists of a pre-trained self-supervised speech model (wav2vec2-large-xlsr-53444https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
@@
We define the logit as the difference between the two similarity scores as
@@
which measures how much more similar the separated signal $\hat{s}_{1}$ is to the text prompt compared to $\hat{s}_{2}$. The logit is scaled by a temperature parameter $T$ ($T<1$) and mapped to a probability using the sigmoid function:
$\hat{y}=\sigma(\text{logit}/T),$
where $\hat{y}\in[0,1]$ represents the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the prompt. A temperature $T<1$ sharpens the output distribution, making predictions more confident (closer to 0 or 1). Empirically, we find that $T=0.2$ yields good performance. The predicted target speaker index is then given by
@@
During inference, the signal with the higher similarity score is selected as the target speech $\hat{s}_{\text{tar}}$.
@@ IV Evaluation
The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems and investigate the relative effectiveness of different cues within the two-stage framework, and (ii) examine whether relative cues yield better performance than independent cues. The following sections provide an overview of the experimental dataset and training configurations, followed by detailed descriptions of each experiment, including the setup, results, and analysis.
@@ IV-A Dataset
To ensure broad linguistic and acoustic diversity, we construct audio mixtures using the same corpora as in our prior work[7]. Each mixture contains speech from two speakers—one designated as the target and the other as the interference—together with a text prompt describing their relative characteristics.
@@
(1) Extraction of Speech Attributes: To provide ground-truth labels for the relative cue categories in each mixture, a set of speech attributes is extracted for every utterance. These attributes serve as the basis for constructing inter-speaker relative cues (e.g., higher pitch, faster speaking rate). All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or reverberation.
@@
Mean $F_{0}$ and $F_{0}$ span are extracted using the Librosa pYIN library [^6] https://librosa.org/doc/main/generated/librosa.pyin.html. When computing speaking duration, word- and pause-level durations are first obtained. If transcription annotations are available, these durations are extracted from time boundaries provided by the Montreal Forced Aligner [^7] https://montreal-forced-aligner.readthedocs.io/en/latest/; otherwise, they are estimated using the Silero voice activity detection tool [^8] https://github.com/snakers4/silero-vad. Speaking duration is then defined as the sum of word durations and short pauses, excluding pauses longer than 0.6 s, which are treated as silence. Speaking rate is computed as the total number of syllables divided by the speaking duration. Syllable counts are estimated based on vowel sequences: in Chinese, each character is treated as one syllable, whereas in English, French, German, and Spanish, syllables are approximated by counting contiguous vowel groups within words. The RMS energy is computed from reverberant speech, averaged over active speech segments to account for speaking duration. Since the mixtures include room acoustic simulations, the speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone. Language, gender, age, emotional state, and transcription are obtained directly from the original dataset annotations.
@@
(2) Audio Mixtures: Each corpus is divided into non-overlapping training, validation, and test subsets, with disjoint speakers and no shared utterances. We generated 100,000 mixtures for training and 10,000 mixtures each for validation and testing. To create realistic speech mixtures, each mixture is composed of two speech signals, $S_{1}$ and $S_{2}$, from different speakers. Both signals are first processed to remove initial and trailing silence, using forced alignment for transcribed speech or voice activity detection for untranscribed segments, and are subsequently temporally arranged with partial overlaps to mimic conversational dynamics. If either $S_{1}$ or $S_{2}$ is shorter than 3 s, the overlap duration is set to the length of the shorter signal, and zero-padding is applied at the beginning of the shorter signal using a randomly selected start offset uniformly drawn from $0,\text{duration of longer sample}-\text{duration of shorter sample}]$. For longer signals, the overlap duration is set to $(\text{length}(S_{1})+\text{length}(S_{2})-6)$ s, with $S_{2}$ padded using a fixed start offset of $6-\text{length}(S_{2})$ s.
@@
To simulate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library[9], representing different speaker positions. Room dimensions are randomly sampled within 9–11 m (length/width) and 2.6–3.5 m (height), with RT60 values of 0.3–0.6 s. The microphone is positioned at the room center, and speakers are placed such that horizontal speaker–microphone distances range from 0.3 to 1.5 m and heights from 1.6 to 1.9 m. The reverberant $S_{2}$ is fixed while $S_{1}$ is scaled to a signal-to-interference ratio (SIR) uniformly sampled between –6 and 6 dB. For each mixture, one of the reverberant signals is randomly selected as the target speech.
@@
(3) Text Prompts: The details of text prompts generation are described in Section II, Table III summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power. Since the “random” and “all” cues are not available in every mixture to specify the target speaker, the number of samples associated with these cues does not correspond to the total number of mixtures.
@@
TABLE III: Sample counts of each relative cue in the training, validation, and test sets, excluding cues labeled as “Similar” or “Same”
@@ IV-B Training Configurations
For the speech encoder in the classification model of the proposed two-stage framework, all cue types span linguistic, paralinguistic, and acoustic information, making the encoder central to capturing these diverse representations. We adopt the encoder from the pretrained wav2vec2-large-xlsr-53 model, a multilingual self-supervised speech representation model based on the wav2vec 2.0 architecture [4]. Trained on speech from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, it learns universal, language-independent representations directly from raw audio. Compared to the ECAPA-TDNN speaker encoder, which focuses on speaker-specific embeddings, wav2vec2-large-xlsr-53 captures richer linguistic, paralinguistic, and acoustic features, making it well suited for cross-lingual target speech extraction. We use the first five encoder layers, with fine-tuning applied to the fifth layer.
@@
To encode text prompts into embeddings, we use the LLaMA 3.2 1B Instruct model for both single-stage TSE systems and the classification model in the two-stage framework. The LLaMA 3.2 1B instruction model is fine-tuned with LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
@@
To ensure high-quality signal separation, we adopt the TF-Locoformer [39], which demonstrates state-of-the-art performance and robustness under reverberation. Single-stage TSE and two-stage separation models are trained with negative SI-SDR loss, float16 mixed precision, and a batch size of 24 (24 V100 GPUs, batch size 1 per GPU). The classification models also use float16 precision and batch size 24.
@@
All models are optimized using AdamW [27], with learning rates and gradient clipping configured individually for each model. The dual-path transformer (DPT)-based single-stage TSE model (used in the following subsection) was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and 1000-step linear warm-up. The TF-Locoformer-based single-stage TSE was trained with a learning rate of 1e-4 and a maximum norm of 5, whereas the two-stage separation model used a learning rate of 1e-3 with the same maximum norm. The classification model use a learning rate of $1\times 10^{-4}$ with max norm 30. Different learning rates and maximum norm values were adopted to account for the distinct architectures and convergence behaviors of the respective models. Learning rates are halved if validation loss does not improve for three epochs, and training employs early stopping with a maximum of 100 epochs.
@@
TABLE IV: Comparison of single-stage and two-stage TSE systems across different cues
@@
IV-C Two-Stage Method versus Single-Stage Baselines
Experimental Setup: We first evaluate the proposed two-stage method against single-stage baselines. In the first stage, the separation model is trained from scratch to obtain separated speech signals. In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, with each prompt corresponding to one of the fourteen relative cue types listed in Table III. The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model, obtained by adapting the TF-Locoformer speech separation model for target speech extraction. Specifically, the model is reconfigured to predict only the target speaker by reducing the number of output speakers in the transposed convolution layer from 2 to 1. Text prompts categorized as “same” or “similar” are excluded from training, validation, and testing, as they provide limited discriminative power.
We also train and evaluate a classification model to test an audio-based TSE approach. Here, the text encoder in the two-stage system is replaced with the audio encoder. For each mixture, a clean utterance from the same target speaker is randomly selected, ensuring that the same utterance is not always used for a given speaker. This setup evaluates the effectiveness of using an enrollment audio cue instead of textual relative cues for target speaker extraction. The clean utterance is normalized to match the mixture length via truncation or zero-padding: signals longer than the target duration are randomly cropped, while shorter signals are zero-padded with randomly distributed offsets. It is then convolved with a RIR randomly drawn from the RIR set described in Section II, producing the reverberant enrollment utterance. To ensure a fair comparison with textual cues, the same number of mixture samples is used for training and validation. Enrollment utterances are distinct from the corresponding mixtures, and evaluation is conducted on mixture sets of the same size as those used for each textual cue, enabling a direct comparison between audio- and text-based TSE approaches.
@@ IV-C Two-Stage Method versus Single-Stage Baselines
Results and Analysis: We evaluate target speech extraction using scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [22] and perceptual evaluation of speech quality (PESQ) [38]. For the TF-Locoformer-based two-stage framework, we report classification accuracy on separated speech (Sep ACC) and overall TSE performance (SI-SDRi/PESQ), computed as the mean score between predicted and ground-truth target speech under different relative cue conditions. Results are summarized in Table IV.
@@
We evaluate three prompt types introduced in Section II (Individual, All, and Random) on their corresponding mixtures. The two-stage framework also supports target speaker enrollment, selecting the separated signal with the highest cosine similarity to the enrollment embedding, enabling direct comparison with text-based cues. TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based counterpart, achieving over 3 dB SI-SDRi and 0.7–1.1 PESQ improvements, except for the age cue, which remains challenging. This suggests that age prediction performance is limited by cue discriminability rather than model architecture. Classification performance across cue types reflects their contribution to single-stage baselines. Compared to single-stage TF-Locoformer-based TSE, the two-stage method shows clear gains: over 3 dB SI-SDRi improvement for random, all, language, transcription, gender, and speaking rate cues, and over 2 dB for most others (except age). It also yields an average PESQ improvement exceeding 0.5, demonstrating strong practical potential.
Some relative cues, such as gender, emotion, and pitch, perform competitively compared with enrollment audio. Interestingly, certain cue types—including all cues, random cues, transcription, loudness, distance, temporal order, and speaking duration—can even outperform the target speaker’s enrollment audio (see Table V, Sep ACC: textual cues vs. enrollment audio). This suggests that enrollment speech may contain confounding information, as it could be recorded in a different environment or exhibit speech characteristics differing from the target speech. These findings highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging discriminative cue types present in the mixture.
@@
To establish an oracle benchmark, we train separate classification models on reverberant–clean speech pairs (Section II, Audio Mixtures of Dataset) using both relative cues and enrollment utterances. We use identical prompt–sample pairs as in the separated-speech setting. The evaluation results are reported in Table V.
TABLE V: Performance of classification across different cues
Cue Type Textual Cues Enrollment Audio
Clean ACC Sep ACC Clean ACC Sep ACC
Random Cues 99.3% 99.2% 95.7% 95.7%
All Cues 99.8% 99.8% 95.7% 95.7%
Language 99.5% 99.3% 96.6% 96.6%
Transcription 90.8% 90.4% 95.8% 95.9%
Gender 98.0% 98.1% 98.2% 98.1%
Emotion 85.8% 87.2% 88.4% 88.6%
Pitch 95.2% 94.6% 95.8% 95.8%
Pitch Range 88.4% 88.7% 95.9% 96.0%
Loudness 98.8% 97.2% 95.7% 95.7%
Distance 99.0% 98.4% 96.2% 96.4%
Age 64.0% 64.3% 93.4% 93.4%
Temporal Order 100% 99.8% 95.6% 95.6%
Speaking Rate 91.5% 91.3% 95.1% 95.4%
Speaking Duration 96.3% 95.9% 95.7% 95.6%
Average 95.8% 95.6% 95.8% 95.8%
Interestingly, the accuracy for each cue type (relative cues and enrollment audio) closely matches that of the model trained on speech separated by the TF-Locoformer model, demonstrating the high quality of the separated signals. The enrollment audio cue shows slightly more consistent classification performance between clean and separated speech, indicating that speaker identity information is more robust than that conveyed by some relative cues, as expected. The variation in classification accuracy across cue types also provides insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
Figure 4: classification accuracy as a function of the attribute difference between target and interfering speakers for each cue type.
We further analyze classification accuracy as a function of the continuous-valued attribute difference between paired target–interference clean speech samples for each cue type. The temporal order cue is excluded due to its perfect classification accuracy (100%). For the remaining cues, logistic regression is used to obtain a smooth estimate of overall classification accuracy over the continuous attribute space. As shown in Fig. 4, accuracy exhibits a clear increasing trend as attribute differences grow. Except for age, all cues reach near-perfect classification when the differences become sufficiently large.